OpenStudy (carlyleukhardt):

What is the value of –1.9 – q when q = 6.7?

Vocaloid (vocaloid):

q = 6.7 means you can replace q with 6.7 whenever you see it

OpenStudy (carlyleukhardt):

oh

Vocaloid (vocaloid):

so -1.9 - q = -1.9 - 6.7 = ?

OpenStudy (carlyleukhardt):

-4.8 ?

Vocaloid (vocaloid):

nope, check your math again

Vocaloid (vocaloid):

-1.9 - 6.7 = -(1.9+6.7) = ?

OpenStudy (carlyleukhardt):

8.6

jimthompson5910 (jim_thompson5910):

as Vocaloid is showing, add `1.9+6.7` then make the result negative

OpenStudy (carlyleukhardt):

-8.6 ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (carlyleukhardt):

thank you :D